OpenStudy (anonymous):

solve for x 6-3(2x-4)=2(1+x)

OpenStudy (anonymous):

Solve for x: 6 x-3 (2 x-4) = 2 (x+1) Factor the polynomial, 2 x-4. Factor 2 from the polynomial 2 x-4: 6 x-3×2 (x-2) = 2 (x+1) Multiply -3 and 2 together. -3×2 = -6: 6 x+-6 (x-2) = 2 (x+1) Distribute -6 over x-2. -6 (x-2) = 12-6 x: 6 x+12-6 x = 2 (x+1) Group like terms in 6 x+12-6 x. Grouping like terms, 6 x+12-6 x = 12+(6 x-6 x): 12+(6 x-6 x) = 2 (x+1) Look for two terms that sum to zero. 6 x-6 x = 0: 12 = 2 (x+1) Reverse the equality in 12 = 2 (x+1) in order to isolate x to the left hand side. 12 = 2 (x+1) is equivalent to 2 (x+1) = 12: 2 (x+1) = 12 Divide both sides by a constant to simplify the equation. Divide both sides of 2 (x+1) = 12 by 2: (2 (x+1))/2 = 12/2 Any nonzero number divided by itself is one. 2/2 = 1: x+1 = 12/2 Reduce 12/2 to lowest terms. The gcd of 12 and 2 is 2, so 12/2 = (2×6)/(2×1) = 2/2×6 = 6: x+1 = 6 Isolate terms with x to the left hand side. Subtract 1 from both sides: x+(1-1) = 6-1 Look for two terms that sum to zero. 1-1 = 0: x = 6-1 Evaluate 6-1. 6-1 = 5:

OpenStudy (anonymous):

x6-3*(2*x-4)=2*(1+x) simplifies to x6-8*x+10=0